First meeting RTGF (Red Temática de Geometría y Física )

Miraflores de la Sierra, Madrid 24–28 November 2003

Organizing Committee: Oscar García-Prada (CSIC, Madrid) César Gómez (IFT, CSIC-UAM, Madrid) Tomás L. Gómez (UCM, Madrid) Angel Uranga (IFT, CSIC-UAM, Madrid)

Programme

MONDAY, 24 November

17h00 – 18h00 Peter Newstead (University of Liverpool and CSIC, Madrid): Moduli spaces of coherent systems

18h30 – 19h30 Luis J. Boya (Universidad de Zaragoza): Euler multiplets, projective geometry and supersymmetry

TUESDAY, 25 November

10h00 – 11h00 Nigel Hitchin (University of Oxford): Geometry with B-fields (I)

11h30 – 12h30 Christoph Schweigert (University of Hamburg): Algebras in tensor categories and conformal field theory (I)

16h00 – 17h00 Ignasi Mundet i Riera (Universidad de Barcelona): Representations of the fundamental group of a Riemann surface in Sp(2n, R)

17h30 – 18h30 Juan J. Manjarin (IFT, UAM-CSIC, Madrid): K-theory and a gravitational Nahm construction WEDNESDAY, 26 November:

10h00 – 11h00 Nigel Hitchin (University of Oxford): Geometry with B-fields (II)

11h30 – 12h30 Christoph Schweigert (University of Hamburg): Algebras in tensor categories and conformal field theory (II)

16h00 – 17h00 Luis Álvarez-Cónsul (CSIC, Madrid): Moduli of quiver bundles

17h30 – 18h30 Adil Belhaj (IFT-UAM, Madrid): Geometric Engineering of N=2 QFT from hyperbolic singularities

THURSDAY, 27 November:

10h00 – 11h00 Richard Thomas (Imperial College, London): Constant scalar curvature Kähler metrics and stability of algebraic varieties

11h30 – 12h30 Vicente Muñoz (UAM,Madrid): Formality of symplectic manifolds

16h00 – 17h00 Jaume Amorós (Universidad Politécnica de Cataluña, Barcelona): Homotopy properties of aspherical symplectic 4-folds

17h30 – 18h30 Luis Fernández (Universidad de los Andes, Bogotá): The dimension of the space of minimal 2-spheres in S^{6}.

FRIDAY, 28 November:

10h00 – 11h00 Richard Thomas (Imperial College, London): Stability of algebraic varieties

11h30 – 12h30 Ruben Minasian (École Polytéchnique, Palaiseau): Mirror symmetric SU(3) manifolds